the demand and supply functions for free range chicken are Pd=50-0.6Q and Ps=20+0.4Q

where P is the price and Q the Quantity, the Equilibrium price and quantity are


1
Expert's answer
2021-05-31T01:58:43-0400

"P_d=P_s"

"50-0.6Q=20+0.4Q"

Equilibrium quantity:

"Q=30"

Equilibrium price:

"P=50-0.6\cdot30=32"